# tablesection algorithmsection

### v1

@@ 3.2 Asymmetric case
×$\hskip 28.45274pt\times\bigg{(}\|\Sigma^{(\pi(2\ell-1),\pi(2\ell-1))}_{Y}\|^{1/2}\|\Sigma^{(\pi(2\ell),\pi(2\ell))}_{Y}\|^{1/2}\varepsilon^{(\pi(2\ell-1),\pi(2\ell))}\bigg{)}$
×$\hskip 28.45274pt\times\bigg{(}\prod_{s=\ell+1}^{p/2}\|\Sigma^{(\pi(2s-1),\pi(2s-1))}_{Y}\|^{1/2}\|\Sigma^{(\pi(2s),\pi(2s))}_{Y}\|^{1/2}\bigg{)}\Bigg{]}$
@@
The bound under the entrywise maximum norm follows from an analogous argument, replacing the supremum over the unit sphere $\mathcal{S}^{d}$ with the supremum over the standard basis =$\mathcal{E}=\{e_{i}\}_{i=1}^{d}$. This completes the proof. ∎
@@ 5 Entrywise maximum norm bound for sample moment tensor
$=\mathbb{E}\max_{v_{1},\dots,v_{p}\in\mathcal{E}}\bigg{|}\frac{1}{N}\sum_{i=1}^{N}\prod_{k=1}^{p}\langle X^{(k)}_{i},v_{k}\rangle-\mathbb{E}\prod_{k=1}^{p}\langle X^{(k)},v_{k}\rangle\bigg{|},$
where $\mathcal{E}$ denotes the standard basis of $\mathbb{R}^{d}$. Let $\bar{\mathcal{E}}:=\mathcal{E}\cup-\mathcal{E}$, and define $\mathcal{F}^{(1)}=\cdots=\mathcal{F}^{(p)}:=\{\langle\cdot,v\rangle:v\in\bar{\mathcal{E}}\}$. Then the maximum can be upper bounded by
@@
$d_{\psi_{2}}(\mathcal{F}^{(k)})=\sup_{f^{(k)}\in\mathcal{F}^{(k)}}\|f^{(k)}\|_{\psi_{2}}\asymp\sup_{f^{(k)}\in\mathcal{F}^{(k)}}\|f^{(k)}\|_{L_{2}}=\sup_{v\in\bar{\mathcal{E}}}\big{(}\mathbb{E}\langle X^{(k)},v\rangle^{2}\big{)}^{1/2}=\|\Sigma^{(k)}\|_{\max}^{1/2}.$
For the generic chaining term, note that the canonical metric on $\bar{\mathcal{E}}$ is given by
@@
$\gamma(\mathcal{F}^{(k)},\psi_{2})\asymp\gamma(\mathcal{F}^{(k)},L_{2})=\gamma(\bar{\mathcal{E}},d_{X^{(k)}})\asymp\mathbb{E}\sup_{u\in\bar{\mathcal{E}}}\langle X^{(k)},u\rangle=\mathbb{E}\|X^{(k)}\|_{\infty}.$

### v2

@@ 3.2 Asymmetric case
$\quad\times\bigg(\|\Sigma^{(\pi(2\ell-1),\pi(2\ell-1))}_{Y}\|^{1/2}\|\Sigma^{(\pi(2\ell),\pi(2\ell))}_{Y}\|^{1/2}\varepsilon^{(\pi(2\ell-1),\pi(2\ell))}\bigg)$
$\quad\times\bigg(\prod_{s=\ell+1}^{p/2}\|\Sigma^{(\pi(2s-1),\pi(2s-1))}_{Y}\|^{1/2}\|\Sigma^{(\pi(2s),\pi(2s))}_{Y}\|^{1/2}\bigg)\Bigg]$
@@
The bound under the entrywise maximum norm follows from an analogous argument, replacing the supremum over the unit spheres $\mathcal{S}^{d_{1}},\ldots,\mathcal{S}^{d_{p}}$ with the supremum over the standard bases $\mathcal{E}^{d_{k}}=\{e_{i}\}_{i=1}^{d_{k}}$ for $1\leq k\leq p$. This completes the proof. ∎
@@ 6 Entrywise maximum norm bound for sample moment tensor
$=\mathbb{E}\max_{v_{k}\in\mathcal{E}^{d_{k}},1\leq k\leq p}\bigg|\frac{1}{N}\sum_{i=1}^{N}\prod_{k=1}^{p}\langle X^{(k)}_{i},v_{k}\rangle-\mathbb{E}\prod_{k=1}^{p}\langle X^{(k)},v_{k}\rangle\bigg|,$
where $\mathcal{E}^{d_{k}}=\{e_{i}\}_{i=1}^{d_{k}}$ denotes the standard basis of $\mathbb{R}^{d_{k}}$. Let $\bar{\mathcal{E}}^{d_{k}}:=\mathcal{E}^{d_{k}}\cup-\mathcal{E}^{d_{k}}$, and define $\mathcal{F}^{(k)}:=\{\langle\cdot,v\rangle:v\in\bar{\mathcal{E}}^{d_{k}}\}$ for $1\leq k\leq p$. Then the maximum can be upper bounded by
@@
$d_{\psi_{2}}(\mathcal{F}^{(k)})=\sup_{f^{(k)}\in\mathcal{F}^{(k)}}\|f^{(k)}\|_{\psi_{2}}\asymp\sup_{f^{(k)}\in\mathcal{F}^{(k)}}\|f^{(k)}\|_{L_{2}}=\sup_{v\in\bar{\mathcal{E}}^{d_{k}}}\big(\mathbb{E}\langle X^{(k)},v\rangle^{2}\big)^{1/2}=\|\Sigma^{(k)}\|_{\max}^{1/2}.$
For the generic chaining term, note that the canonical metric on $\bar{\mathcal{E}}^{d_{k}}$ is given by
@@
$\gamma(\mathcal{F}^{(k)},\psi_{2})\asymp\gamma(\mathcal{F}^{(k)},L_{2})=\gamma(\bar{\mathcal{E}}^{d_{k}},d_{X^{(k)}})\asymp\mathbb{E}\sup_{u\in\bar{\mathcal{E}}^{d_{k}}}\langle X^{(k)},u\rangle=\mathbb{E}\|X^{(k)}\|_{\infty}.$